OpenStudy (anonymous):

Which of the following is the equation of a circle with center (5, -2) and radius of 6?

OpenStudy (anonymous):

a) (x - 5)2 + (y + 2)2 = 36 b) (x - 5)2 + (y + 2)2 = 6 c) (x + 5)2 + (y - 2)2 = 6 d) (x + 5)2 + (y - 2)2 = 36

OpenStudy (cwrw238):

the general equation of a circle can be written (x - a)^2 + (y - b)^2 = r^2 where (a,b) = center and r = radius

OpenStudy (cwrw238):

so if you compare each of your options with this equation you should see which one is correct

OpenStudy (cwrw238):

note if r = 6 then r^2 = 36

OpenStudy (anonymous):

i think it's a)

OpenStudy (cwrw238):

right (x - 5)^2 ( ( x - (-2)) = 6^2

OpenStudy (anonymous):

oh okay i get it now!! thank you for helping! uvu

OpenStudy (cwrw238):

(x - 5)^2 + ( x - (-2)) = 6^2

OpenStudy (cwrw238):

yw